Expected Standard Deviation Return Stock fund (S) Bond fund (B) 17% 30% 22 11 The correlation between the fund returns is 0.1A pension fund manager is considering three mutual funds. The first is a stock fund, the second is a long-term government and corporate bond fund, and the third is a T-bill money market fund that yields a rate of 8%. The probability distribution of the risky funds is as follows:
Answer #1

Given information:

stock fund return rs = 0.17

stock fund standard deviation \sigma_s = 0.3

Bond fund return rb = 0.11

Bond fund standard deviation \sigma_b = 0.22

correlation \rho = 0.1

Part a-1:

Find the investment proportions (weights) of the risky funds in the portfolio.

Assume "w" is the stock fund weight in the portfolio. Since the sum of weights =1, the bond fund weight will be (1-w)

Expected return of the portfolio = weight of stock fund * returns of stock fund + weight of bond fund * returns of bond fund

E (portfolio) = w * 0.17 + (1-w) *0.11 = 0.06 w +0.11

Variance (portfolio) = w2 * \sigma_s 2 + (1-w)2 * \sigma_b 2 + 2 *  \rho * w * (1-w) * \sigma_s * \sigma_b

Substitute the given values:

Variance = w2 * 0.32 + (1-w)2 * 0.222 + 2 * 0.1 * w * (1-w) * 0.3 * 0.22

= 0.09 * w2 + (1-2w+w2 ) *0.0484 +0.0132 (w - w2 )

= 0.1252 * w2 - 0.0836*w +0.0484

In order to find minimum variance portfolio, we will use calculus concepts here. determine the derivative of variance V with respect to weight w and equate it to 0.

dV/dw = 0.1252 * 2 w - 0.0836 = 0

w = 0.0836 / (0.1252*2) = 0.3339

Now take the second derivative and if this is positive, then from the rules of basic calculus, the variance is at its minimum.

d2v/dw2 = 0.1252 * 2 dw/dw = 0.1252 * 2 = 0.2504 > 0

Since the second derivative is positive, the variance is at its minimum.

Answer:

Portfolio invested in stock = 0.3339

Portfolio invested in bond = 1- 0.3339 = 0.6661

Part a-2

Substitute the value of w in the expected return and standard deviation formula

E (portfolio) = w * 0.17 + (1-w) *0.11 = 0.06 w +0.11 =0.06 * 0.3339 + 0.11 = 0.13 = 13%

Variance = 0.1252 * w2 - 0.0836*w +0.0484 = = 0.1252 * 0.33392 - 0.0836*0.3339 +0.0484 = 0.0344

Standard deviation = square root (0.0344) = 0.1856

Answer:

Expected return = 13%

Standard deviation = 18.56%